# The spectral continuum in the Rabi-Stark model

### v1

@@ 1 Introduction
The Rabi model enjoys a long and distinguished career. Originally introduced by Rabi [1] in 1936 as a semi-classical model to describe the interaction of quantum mechanical spins with a classical electromagnetic field, its fully quantum mechanical version was formulated by Jaynes and Cummings [2] in 1963. This model, the quantum Rabi model (QRM), captures the interaction of matter (Fermionic discrete degrees of freedom) and light (Bosonic continuous degrees of freedom) in a minimalist way by assuming that a matter qubit, representing a two-level atom, couples to a single-mode Bosonic field. Jaynes and Cummings simplified the QRM using the so-called rotating-wave approximation (RWA), valid for small matter-field interaction, to obtain a model now bearing the authors’ name, the quantum Jaynes-Cummings model, which they then solved by elementary means [2]. The rotating-wave approximation enhances the $\mathbb{Z}_{2}$-symmetry of the QRM to a continuous $U(1)$-symmetry which renders it solvable in closed form by dividing the Hilbert space into infinitely many two-dimensional invariant subspaces. A similar symmetry enhancement, adapted to the case of strong coupling and called the generalized rotating wave approximation (GRWA), led again to closed form expressions for eigenenergies and eigenstates, overcoming the failure of the RWA for strong coupling [3]. It was shown later that the original $\mathbb{Z}_{2}$-symmetry of the QRM suffices to obtain a closed formula for the spectral determinant of the Hamiltonian, the so-called $G$-function [4]. This symmetry divides the Hilbert space into just two invariant subspaces, each infinite dimensional, such that level crossings between different subspaces become possible. However, these exact crossings are far less numerous than in the approximations with enhanced symmetry, RWA and GRWA. The solution of the QRM given in [4] has been reproduced using Bogoliubov operators [5] by employing a heuristic argument for the implementation of the spectral condition (normalizability of the eigenfunctions).
@@
While the quantum Jaynes-Cummings model was sufficient to interpret experiments for a long time, recent experimental developments have been able to drive the matter-field interaction strength into regimes -- the strong, ultra-strong [6] and deep strong [7] coupling regimes -- where the quantum Jaynes-Cummings model is no longer appropriate, requiring the full QRM as a replacement. Forays into these coupling regimes are an active current research area, with examples ranging from trapped ion systems interacting with a laser field confined to a cavity [8] to superconducting charge qubits in circuit quantum electrodynamics architectures [9, 10].
@@
A remarkable property of the QRM is that it exhibits a “quantum phase transition”, showing universal dynamics in the vicinity of criticality [11]. The appearance of a phase transition in a few-body system comes as a surprise as such a transition may occur only in the thermodynamic limit of a many-body system, when the particle number $N$ and the volume $V$ diverge such that the density $N/V$ stays finite [12]. The thermodynamic limit entails that the level distance between many-body eigenstates goes to zero, a fact which can be mimicked in the QRM by the limit $\Delta/\omega\rightarrow\infty$, $g/\omega\rightarrow\infty$ such that $g/{\sqrt{\omega\Delta}}$ stays finite. The infinite dimensional Hilbert space of the single bosonic mode may replace thus the state space of infinitely many particles, at the cost of introducing experimentally inaccessible parameters. Moreover, this few-body phase transition is of mean-field character [13].
@@
There are many generalizations of the QRM, inter alia the two-qubit QRM [14, 15], the two-photon QRM [16, 17], and the asymmetric QRM [4]. The asymmetric version of the model is particularly intriguing and is currently attracting much attention. While the additional driving term breaks the $\mathbb{Z}_{2}$-invariance and generically lifts the level crossings of the QRM, the crossings reappear for certain discrete values of the amplitude of the driving term [4, 18]. This observation has led to the search for hidden symmetries in the asymmetric QRM and the construction of corresponding operators [19, 20, 21, 22, 23, 24], in addition to an analysis of entanglement resonances as a symmetry probe [25].
@@
A review of the QRM and several of its generalizations is given in [26]. Recent reviews of light-matter interaction models and methods for their solution can be found in [27] and, with an emphasis on the QRM, in [28]. For a recent treatment of the derivation from first principles and an elementary discussion of the QRM, see [29].
@@
Some ten years ago Grimsmo and Parkins proposed yet another generalization of the QRM [30] which will be the model further investigated in the present work. These authors suggested an experimental set-up singling out two hyperfine ground states of a multilevel atom constituting an effective two-level quantum system and coupled to a quantized cavity field (single-frequency photons) and two auxiliary laser fields. The quantum Rabi model realized in this way permits the free and independent choice of frequencies and coupling constants of the model such that also the ultrastrong and deep-strong coupling regimes can be accessed experimentally. Unless the two laser frequencies are fine-tuned, i.e. for generic values of the model parameters, an additional, nonlinear, term appears in the model Hamiltonian coupling the two-level system and the cavity field. Such a term is well known as dynamical Stark shift, a quantum analog of the Bloch-Siegert shift [31]. The model is meanwhile known under the name of quantum Rabi-Stark model (QRSM). In the dynamical Stark shift, the nonlinear coupling strength is usually dependent on the other system parameters of the model while in the Grimso-Parkins scheme also the Stark coupling can be adjusted freely and independently. The emergence of selective $k$-photon interactions in the strong and ultra-strong regimes and an alternative method for the quantum simulation of the QRSM was recently demonstrated [32]. The latter uses the internal and vibrational degrees of freedom of a trapped ion set-up.
@@
Grimsmo and Parkins [33] discuss the possibility that the QRSM can exhibit a superradiant phase transition in the deep-strong coupling regime when the magnitude of the Stark coupling becomes equal to the frequency of the cavity field, i.e., at finite values of the parameters, thus avoiding the unphysical parameter regime which led to the phase transition in the QRM [11]. Spurred by this conjecture, there has been a number of works analyzing the spectral structure of the QRSM [34, 35, 36, 37, 38]. Despite these efforts, a number of intriguing questions remain open, some of which will be addressed in this paper.
@@
and the Hilbert space ${\mathcal{H}}$ decomposes into two subspaces ${\mathcal{H}}={\mathcal{H}}_{+}\oplus{\mathcal{H}}_{-}$. The eigenstates and eigenvalues of this Hamiltonian can thus be labeled by two quantum numbers: the parity quantum number $p=\pm 1$ and an eigenvalue $n=0,1,2,\ldots$ numbering the energy eigenvalues from below. The spectral structure of this Hamiltonian [4] accordingly shows a positive and a negative parity part where level crossings are only possible between energy levels of different parities while levels with the same parity avoid level crossings. The spectrum of (1) can also be described as consisting of a regular part to which almost all eigenvalues belong and an exceptional part formed by the level crossings of opposite parity levels.
The QRSM with model Hamiltonian [^2] We here use the notation from [36] with parameters $\gamma,\Delta$ and $g$ related to the parameters $U,\omega_{0}$, and $g_{\textrm{eff}}$ in [30] by $\gamma=U/2,\Delta=\omega_{0}/2$, and $g=g_{\textrm{eff}}$. Other conventions can be found in the literature, see e.g. [34, 35, 38].
@@
also enjoys the $\mathbb{Z}_{2}$ symmetry and, hence, its spectrum can, on the whole, be described similar to the QRM spectrum. Indeed, we shall show in section 2 that the coupled system of ordinary differential equations (ODE’s) representing the eigenvalue equation for $H_{RS}$ can be mapped onto the corresponding system of ODE’s for $H_{R}$, albeit with certain modifications which result in qualitative changes which we now describe phenomenologically.
The first characteristic of the QRSM is that, whereas in the QRM each successive pair of spectral levels labeled by $\{p=\pm 1,n=0,1,2,\ldots\}$ exhibits $n$ level crossings before it eventually becomes degenerate for Rabi coupling $g\rightarrow\infty$, the QRSM exhibits one more level crossing before the levels again become pairwise degenerate for $g\rightarrow\infty$. The additional level crossings occur for [38]
@@
Note that all $g_{c}^{(n)}\rightarrow\infty$ as $\gamma\rightarrow 0$, reproducing the behaviour of the QRM spectrum. The first of the additional crossings, between the ground state and the first excited state, occurring at $g_{c}^{(0)}$, has been interpreted as a first-order quantum phase transition (QPT) [38] which occurs in similar way in the anisotropic QRM [39].
@@
While the former levels, for $\gamma\rightarrow\omega$, eventually tend to the spectrum of a harmonic oscillator, the latter tend, in the same limit, to a massively degenerate spectrum at $E=-\Delta$. However, as we shall see in the following, there is a dramatic restructuring of the spectrum when switching on $g$: When $\gamma\rightarrow\omega$ the latter branch of energy levels now fuse into a continuum of states similar to the QRM with non-linear (two-photon) coupling when the coupling reaches a critical value [16, 17, 40, 41].
@@
We shall have occasion to interpret the spectral structure of the QRSM in the limiting case $\gamma\rightarrow\omega$ with $g>0$ as spectral levels of bound states embedded in a continuum of non-normalizable states. This interpretation is borne out by the results of large scale numerical exact diagonalization calculations which reveal the phenomenological picture of the spectrum outlined above. Moreover, analytical calculations demonstrate explicitly what the numerical spectra suggest and, furthermore, in the limit $\gamma\rightarrow\omega$, prove the existence of discrete parts of the spectrum corresponding to bound states, and continuous parts of the spectrum corresponding to extended states. The bound states are embedded in the continuum and can thus be identified as “bound states in the continuum” (BICs), a phenomenon which has attracted considerable experimental and theoretical attention in recent years, see the review [42] and the following discussion.
@@
Returning to the original proposal for the QRSM [30], Grimsmo and Parkins noted that an instability develops in the system as the magnitude of the Stark coupling $|\gamma|$ grows towards the photon frequency $\omega$, causing sharp changes in the photon number and other observables in the deep-strong-coupling regime $g>\omega$. It was conjectured that $\gamma=\pm\omega$ define critical points that mark the occurrence of quantum phase transitions, with the ground states here becoming massively degenerate. Specifically -- based on a semiclassical analysis of the reduced density matrix in the presence of cavity dissipation -- the system was predicted to undergo a transition to a superradiant phase at $\gamma=-\omega$ [30].
@@
Independent support for the existence of possible phase transitions at $\gamma=\pm\omega$ was provided by Xie et al., using a mapping of the QRSM Hamiltonian (3) at the critical point $\gamma=\omega$ to an effective quantum harmonic oscillator [38]. These authors found that part of the discrete energy spectrum collapses onto a massively degenerate level at a critical Rabi coupling $g_{c}=\sqrt{(\omega/2-\Delta)\omega}$, exhibiting manifest scaling behavior of the photon number indicative of a quantum phase transition [43].  [^3] We here write the expression for $g_{c}$ with unnormalized $\omega$, while in [30] $\omega$ was normalized to unity. Also, adhering to the notation in [36], our parameter $\Delta$ in the same expression differs by a factor of 2 compared to the corresponding parameter in [30] and [38]. Intriguingly, for all $g>0$, still at $\gamma=\pm\omega$, there is the possibility that the spectral gap for energies $E$ in the interval $-\Delta-2g^{2}/\omega<E<-\Delta$ may get filled by a continuous band of energies. If so, this interval supports an uncountable set of non-normalizable (scattering) states. A similar appearance of non-normalizable states in the two-photon Rabi model is well known [40, 41] and was anticipated also for the QRSM by Maciejewski et al. [35]. However, physical implications -- as well as a rigorous proof of the very existence of a continuum of states -- have remained elusive.
@@
In this article we revisit the problem, taking off from a set of scale transformations that map the Hamiltonian in (3) to an effective Rabi model. Similar to the ordinary Rabi model, the eigenvalue equation becomes equivalent to a coupled system of first-order ODE’s in Bargmann space [4]. The two regular singular points of this system are found to coalesce to an irregular singular point at infinity in the limit $\gamma\rightarrow\omega_{-}$. By studying the confluence process, with the standard assumption that the variable prefactors in the system of ODE’s are bounded with respect to the diverging parameters, we can furnish a formal proof that there is a spectral continuum with non-normalizable states between $-\Delta-2g^{2}/\omega$ and $-\Delta$. For $E\geq-\Delta$ the spectrum remains discrete with normalizable eigenstates.
@@
Lifting the assumption that the confluence process comes with a certain hierarchy of diverging quantities unveils a dramatic scenario: We can now prove the emergence of an unbounded continuous energy band for all energies $E>-\Delta-2g^{2}/\omega$ in the limit $\gamma\rightarrow\omega$. The corresponding non-normalizable (scattering) states are orthogonal to the normalizable states of the discrete spectrum for $E>-\Delta$, implying that the latter states get embedded in a spectral continuum. It follows that the normalizable states can be considered as bound states in the continuum (BICs) [42, 44, 45] at the point $\gamma=\omega$ (with the expectation -- yet to be proven -- that the same scenario plays out also at $\gamma=-\omega$).
@@
The anomalous confinement phenomenon implied by BICs has attracted a lot of interest recently, primarily brought on by potential applications in photonics [46, 47, 48, 49]. Quite generally, the existence of a BIC -- be it in the classical or quantum realm -- depends on some mechanism that confines the state despite it being degenerate with an extended scattering state (that would otherwise radiate away the BIC under an infinitesimal perturbation). Numerous confinement mechanisms have been proposed over the years -- from tailored potentials (oscillating in such a way as to support a BIC, as in the pioneering work by von Neumann and Wigner [50]), to symmetry protection (with no coupling between the BIC and the extended states if belonging to different symmetry classes [51, 52]) and separation of variables (with each term of a separable Hamiltonian having a bounded eigenstate in addition to a continuous band [53, 54]).
@@
While these classes of BICs and their origins have been well studied -- albeit still challenging to realize in a quantum system -- the BICs we find in the QRSM are of a different kind. The only requirement here is the fine tuning that sets $\gamma=\omega$. One may think that this is an instance of a parametric BIC, yet another class of such states (where a BIC gets confined by tuning the radiating channels so that they cancel [55]). However, different from generic parametric BICs known from the literature, only a single parameter needs tuning in our case, making it potentially more accessible in an experiment.
@@
The paper is organized as follows: In section 2, we calculate the spectral determinant ($G$-function) $G_{\pm}(x)$ of the QRSM in a simplified way by mapping it to an effective QRM. In this way, results obtained in [38] about a novel class of degeneracies and the concomitant first order quantum phase transition are easily explained. In section 3, we study the limit $\gamma\rightarrow\omega_{-}$. The pole structure of $G_{\pm}(x)$ gives information about the average level spacing and thus a first hint to the presence of a continuum at $\gamma=\omega$ for $E>-\Delta$. There are two possible confluence procedures: the first one is standard and has been employed in [35, 38]. The second one is novel and yields the full continuum together with the embedded bound states. In section 4 we give independent corroborating evidence for the full continuum by numerical exact diagonalization of the model in state spaces with very high dimension. Section 5 is devoted to an independent confirmation through the explicit analytical treatment of the QRSM under the assumption that the radiation mode is slow compared to the qubit. Finally, we present our conclusions and outlook to future research directions in section 6.
2 Effective quantum Rabi model
The following analysis uses the representation of the creation operator $a^{\dagger}$ as multiplication with the complex variable $z$ and the annihilation operator $a$ as derivative ${\rm{d}}/{\rm{d}}z$. These operators are mutually adjoint in the Bargmann space ${\cal B}$ of analytic functions [56]. States $\psi(x)\in L^{2}(\mathbb{R})$, the Hilbert space of square integrable functions on the real line, are mapped to analytic functions $\psi(z)$ of $z$. The inner product is defined as

### v2

@@ I Introduction
The Rabi model enjoys a long and distinguished career. Originally introduced by Rabi Rabi1936 in 1936 as a semi-classical model to describe the interaction of quantum mechanical spins with a classical electromagnetic field, its fully quantum mechanical version was formulated by Jaynes and Cummings JC in 1963. This model, the quantum Rabi model (QRM), captures the interaction of matter (Fermionic discrete degrees of freedom) and light (Bosonic continuous degrees of freedom) in a minimalist way by assuming that a matter qubit, representing a two-level atom, couples to a single-mode Bosonic field. Jaynes and Cummings simplified the QRM using the so-called rotating-wave approximation (RWA), valid for small matter-field interaction, to obtain a model now bearing the authors’ name, the quantum Jaynes-Cummings model, which they then solved by elementary means JC .
The rotating-wave approximation enhances the $\mathbb{Z}_{2}$-symmetry of the QRM to a continuous $U(1)$-symmetry which renders it solvable in closed form by dividing the Hilbert space into infinitely many two-dimensional invariant subspaces. A similar symmetry enhancement, adapted to the case of strong coupling and called the generalized rotating wave approximation (GRWA), led again to closed form expressions for eigenenergies and eigenstates, overcoming the failure of the RWA for strong coupling irish2007 . It was shown later that the original $\mathbb{Z}_{2}$-symmetry of the QRM suffices to obtain a closed formula for the spectral determinant of the Hamiltonian, the so-called $G$-function braak2011 . This symmetry divides the Hilbert space into just two invariant subspaces, each infinite dimensional, such that level crossings between different subspaces become possible. However, these exact crossings are far less numerous than in the approximations with enhanced symmetry, RWA and GRWA. The solution of the QRM given in braak2011 has been reproduced using Bogoliubov operators chen2012 by employing a heuristic argument for the implementation of the spectral condition (normalizability of the eigenfunctions).
@@
While the quantum Jaynes-Cummings model was sufficient to interpret experiments for a long time, recent experimental developments have been able to drive the matter-field interaction strength into regimes -- the strong, ultra-strong niemczyk2010 and deep strong casanova2010 coupling regimes -- where the quantum Jaynes-Cummings model is no longer appropriate, requiring the full QRM as a replacement. Forays into these coupling regimes are an active current research area, with examples ranging from trapped ion systems interacting with a laser field confined to a cavity haroche2008 to superconducting charge qubits in circuit quantum electrodynamics architectures blais2020 ; blais2021 .
@@
A remarkable property of the QRM is that it exhibits a “quantum phase transition”, showing universal dynamics in the vicinity of criticality bakemeier2012 ; ashhab2013 ; hwang2015 . The appearance of a phase transition in a few-body system comes as a surprise as such a transition may occur only in the thermodynamic limit of a many-body system, when the particle number $N$ and the volume $V$ diverge such that the density $N/V$ stays finite sachdev2011 . The thermodynamic limit entails that the level distance between many-body eigenstates goes to zero, a fact which can be mimicked in the QRM by the limit $\Delta/\omega\rightarrow\infty$, $g/\omega\rightarrow\infty$ such that $g/{\sqrt{\omega\Delta}}$ stays finite. The infinite dimensional Hilbert space of the single bosonic mode may replace thus the state space of infinitely many particles, at the cost of introducing experimentally inaccessible parameters. Moreover, this few-body phase transition is of mean-field character larson2017 .
@@
There are many generalizations of the QRM, inter alia the two-qubit QRM peng2012 ; sun2020 , the two-photon QRM duan2016 ; cong2019 ; rico2020 ; braak2023 , and the asymmetric QRM braak2011 . The asymmetric version of the model is particularly intriguing and is currently attracting much attention. While the additional driving term breaks the $\mathbb{Z}_{2}$-invariance and generically lifts the level crossings of the QRM, the crossings reappear for certain discrete values of the amplitude of the driving term braak2011 ; ashhab2020 . This observation has led to the search for hidden symmetries in the asymmetric QRM and the construction of corresponding operators mangazeev2021 ; li2021a ; Kimoto2021 ; reyes2021 ; li2021b ; xie2021 , in addition to an analysis of entanglement resonances as a symmetry probe shi2022 .
@@
A review of the QRM and several of its generalizations is given in xie2017 . Recent reviews of light-matter interaction models and methods for their solution can be found in leboite2020 and, with an emphasis on the QRM, in larson2021 . For a recent treatment of the derivation from first principles and an elementary discussion of the QRM, see eckle2019 .
@@
Some ten years ago Grimsmo and Parkins proposed yet another generalization of the QRM grimsmo2013 which will be the model further investigated in the present work. These authors suggested an experimental set-up singling out two hyperfine ground states of a multilevel atom constituting an effective two-level quantum system and coupled to a quantized cavity field (single-frequency photons) and two auxiliary laser fields. The quantum Rabi model realized in this way permits the free and independent choice of frequencies and coupling constants of the model such that also the ultrastrong and deep-strong coupling regimes can be accessed experimentally. Unless the two laser frequencies are fine-tuned, i.e. for generic values of the model parameters, an additional, nonlinear, term appears in the model Hamiltonian coupling the two-level system and the cavity field. Such a term is well known as dynamical Stark shift, a quantum analog of the Bloch-Siegert shift klimov2009 . The model is meanwhile known under the name of quantum Rabi-Stark model (QRSM). In the dynamical Stark shift, the nonlinear coupling strength is usually dependent on the other system parameters of the model while in the Grimso-Parkins scheme also the Stark coupling can be adjusted freely and independently. The emergence of selective $k$-photon interactions in the strong and ultra-strong regimes and an alternative method for the quantum simulation of the QRSM was recently demonstrated cong2020 . The latter uses the internal and vibrational degrees of freedom of a trapped ion set-up.
@@
Grimsmo and Parkins grimsmo2014 discuss the possibility that the QRSM can exhibit a superradiant phase transition in the deep-strong coupling regime when the magnitude of the Stark coupling becomes equal to the frequency of the cavity field, i.e., at finite values of the parameters, thus avoiding the unphysical parameter regime which led to the phase transition in the QRM hwang2015 . Spurred by this conjecture, there has been a number of works analyzing the spectral structure of the QRSM maciejewski2014 ; maciejewski2015 ; eckle2017 ; eckle2023 ; xie2019 . Despite these efforts, a number of intriguing questions remain open, some of which will be addressed in this paper.
@@
and the Hilbert space ${\mathcal{H}}$ decomposes into two subspaces ${\mathcal{H}}={\mathcal{H}}_{+}\oplus{\mathcal{H}}_{-}$. The eigenstates and eigenvalues of this Hamiltonian can thus be labeled by two quantum numbers: the parity quantum number $p=\pm 1$ and an eigenvalue $n=0,1,2,\ldots$ numbering the energy eigenvalues from below. The spectral structure of this Hamiltonian braak2011 accordingly shows a positive and a negative parity part where level crossings are only possible between energy levels of different parities while levels with the same parity avoid level crossings. The spectrum of (1) can also be described as consisting of a regular part to which almost all eigenvalues belong and an exceptional part formed by the level crossings of opposite parity levels.
The QRSM with model Hamiltoniannote1
@@
also enjoys the $\mathbb{Z}_{2}$ symmetry and, hence, its spectrum can, on the whole, be described similar to the QRM spectrum. Indeed, we shall show in section II that the coupled system of ordinary differential equations (ODE’s) representing the eigenvalue equation for $H_{RS}$ can be mapped onto the corresponding system of ODE’s for $H_{R}$, albeit with certain modifications which result in qualitative changes which we now describe phenomenologically.
The first characteristic of the QRSM is that, whereas in the QRM each successive pair of spectral levels labeled by $\{p=\pm 1,n=0,1,2,\ldots\}$ exhibits $n$ level crossings before it eventually becomes degenerate for Rabi coupling $g\rightarrow\infty$, the QRSM exhibits one more level crossing before the levels again become pairwise degenerate for $g\rightarrow\infty$. The additional level crossings occur for xie2019
@@
Note that all $g_{c}^{(n)}\rightarrow\infty$ as $\gamma\rightarrow 0$, reproducing the behaviour of the QRM spectrum. The first of the additional crossings, between the ground state and the first excited state, occurring at $g_{c}^{(0)}$, has been interpreted as a first-order quantum phase transition (QPT) xie2019 which occurs in similar way in the anisotropic QRM chen2021 .
@@
While the former levels, for $\gamma\rightarrow\omega$, eventually tend to the spectrum of a harmonic oscillator, the latter tend, in the same limit, to a massively degenerate spectrum at $E=-\Delta$. However, as we shall see in the following, there is a dramatic restructuring of the spectrum when switching on $g$: When $\gamma\rightarrow\omega$ the latter branch of energy levels now fuse into a continuum of states similar to the QRM with non-linear (two-photon) coupling when the coupling reaches a critical value duan2016 ; rico2020 ; braak2023 ; ng1999 ; felicetti2015 .
@@
We shall have occasion to interpret the spectral structure of the QRSM in the limiting case $\gamma\rightarrow\omega$ with $g>0$ as spectral levels of bound states embedded in a continuum of non-normalizable states. This interpretation is borne out by the results of large scale numerical exact diagonalization calculations which reveal the phenomenological picture of the spectrum outlined above. Moreover, analytical calculations demonstrate explicitly what the numerical spectra suggest and, furthermore, in the limit $\gamma\rightarrow\omega$, prove the existence of discrete parts of the spectrum corresponding to bound states, and continuous parts of the spectrum corresponding to extended states. The bound states are embedded in the continuum and can thus be identified as “bound states in the continuum” (BICs), a phenomenon which has attracted considerable experimental and theoretical attention in recent years, see the review BICreview and the following discussion.
@@
Returning to the original proposal for the QRSM grimsmo2013 , Grimsmo and Parkins noted that an instability develops in the system as the magnitude of the Stark coupling $|\gamma|$ grows towards the photon frequency $\omega$, causing sharp changes in the photon number and other observables in the deep-strong-coupling regime $g>\omega$. It was conjectured that $\gamma=\pm\omega$ define critical points that mark the occurrence of quantum phase transitions, with the ground states here becoming massively degenerate. Specifically -- based on a semiclassical analysis of the reduced density matrix in the presence of cavity dissipation -- the system was predicted to undergo a transition to a superradiant phase at $\gamma=-\omega$ grimsmo2013 .
@@
Independent support for the existence of possible phase transitions at $\gamma=\pm\omega$ was provided by Xie et al., using a mapping of the QRSM Hamiltonian (3) at the critical point $\gamma=\omega$ to an effective quantum harmonic oscillator xie2019 . These authors found that part of the discrete energy spectrum collapses onto a massively degenerate level at a critical Rabi coupling $g_{c}=\sqrt{(\omega/2-\Delta)\omega}$, exhibiting manifest scaling behavior of the photon number indicative of a quantum phase transition chen2020 ; note2 Intriguingly, for all $g>0$, still at $\gamma=\pm\omega$, there is the possibility that the spectral gap for energies $E$ in the interval $-\Delta-2g^{2}/\omega<E<-\Delta$ may get filled by a continuous band of energies. If so, this interval supports an uncountable set of non-normalizable (scattering) states. A similar appearance of non-normalizable states in the two-photon Rabi model is well known rico2020 ; braak2023 and was anticipated also for the QRSM by Maciejewski et al. maciejewski2015 . However, physical implications -- as well as a rigorous proof of the very existence of a continuum of states -- have remained elusive.
@@
In this article we revisit the problem, taking off from a set of scale transformations that map the Hamiltonian in Equation (3) to an effective quantum Rabi model. Similar to the ordinary QRM, the eigenvalue equation becomes equivalent to a coupled system of first-order ODE’s in Bargmann space braak2011 . The two regular singular points of this system are found to coalesce to an irregular singular point at infinity in the limit $\gamma\rightarrow\omega_{-}$. By studying the confluence process, with the standard assumption that the variable prefactors in the system of ODE’s are bounded with respect to the diverging parameters, we can furnish a formal proof that there is a spectral continuum with non-normalizable states between $-\Delta-2g^{2}/\omega$ and $-\Delta$. For $E>-\Delta$ the spectrum remains discrete with normalizable eigenstates.
@@
Lifting the assumption that the confluence process comes with a certain hierarchy of diverging quantities unveils a dramatic scenario: We can now prove the emergence of an unbounded continuous energy band for all energies $E>-\Delta-2g^{2}/\omega$ in the limit $\gamma\rightarrow\omega$. The corresponding non-normalizable (scattering) states are orthogonal to the normalizable states of the discrete spectrum for $E>-\Delta$, implying that the latter states get embedded in a spectral continuum. It follows that the normalizable states can be considered as bound states in the continuum (BICs) BICreview ; koshelev2021 ; sadreev2021 at the point $\gamma=\omega$ (with the expectation -- yet to be proven -- that the same scenario plays out also at $\gamma=-\omega$).
@@
The anomalous confinement phenomenon implied by BICs has attracted a lot of interest recently, primarily brought on by potential applications in photonics kodigala2017 ; longhi2021 ; qin2022 ; kang2023 . Quite generally, the existence of a BIC -- be it in the classical or quantum realm -- depends on some mechanism that confines the state despite it being degenerate with an extended scattering state (that would otherwise radiate away the BIC under an infinitesimal perturbation). Numerous confinement mechanisms have been proposed over the years -- from tailored potentials (oscillating in such a way as to support a BIC, as in the pioneering work by von Neumann and Wigner neumann1929 ), to symmetry protection (with no coupling between the BIC and the extended states if belonging to different symmetry classes plotnik2011 ; zhang2012 ) and separation of variables (with each term of a separable Hamiltonian having a bounded eigenstate in addition to a continuous band jain1975 ; robnik1986 ).
@@
While these classes of BICs and their origins have been well studied -- albeit still challenging to realize in a quantum system -- the BICs we find in the QRSM are of a different kind. The only requirement here is the fine tuning that sets $\gamma=\omega$. One may think that this is an instance of a parametric BIC, yet another class of such states (where a BIC gets confined by tuning the radiating channels so that they cancel hsu2013 ). However, different from generic parametric BICs known from the literature, only a single parameter needs tuning in our case, making it potentially more accessible in an experiment.
@@
The paper is organized as follows: In section II, we calculate the spectral determinant ($G$-function) $G_{\pm}(x)$ of the QRSM in a simplified way by mapping it to an effective QRM. In this way, results obtained in xie2019 about a novel class of degeneracies and the concomitant first order quantum phase transition are easily explained. In section III, we study the limit $\gamma\rightarrow\omega_{-}$. The pole structure of $G_{\pm}(x)$ gives information about the average level spacing and thus a first hint to the presence of a continuum at $\gamma=\omega$ for $E>-\Delta$. There are two possible confluence procedures: the first one is standard and has been employed in maciejewski2015 ; xie2019 . The second one is novel and yields the full continuum together with the embedded bound states. In section IV we give independent corroborating evidence for the full continuum by numerical exact diagonalization of the model in state spaces with very high dimension. Section V is devoted to an independent confirmation through the explicit analytical treatment of the QRSM under the assumption that the radiation mode is slow compared to the qubit. Finally, we present our conclusions and outlook to future research directions in section VI.
II Effective quantum Rabi model
The following analysis uses the representation of the creation operator $a^{\dagger}$ as multiplication with the complex variable $z$ and the annihilation operator $a$ as derivative ${\rm{d}}/{\rm{d}}z$. These operators are mutually adjoint in the Bargmann space ${\cal B}$ of analytic functions bargmann1961 . States $\psi(x)\in L^{2}(\mathbb{R})$, the Hilbert space of square integrable functions on the real line, are mapped to analytic functions $\psi(z)$ of $z$. The inner product is defined as